Required information The following information applies to the questions displayed below. Data for Hermann Corporation are shown below Percent Per Unit of Sales 108% Selling price var1able expenses Contribution margin S 90 63 ed S 27 30% Fixed expenses are $30,000 per month and the company is selling 2.000 units per month. 2-o. Refer to the original data. How much will net operating income increase (decrease) per month if the company uses higher-quality components that increase the variable expense by $2 per unit and increase unit sales by 10%. 2-b. Should the higher-quality components be used? Complete this question by entering your answers in the tabs below Req 2A Req 2B Refer to the original data. How much will net operating income increase (decrease) per month if the company uses higher- quality components that increase the variable expense by $2 per unit and increase unit sales by 10% decreases 8 Req 2A Req 2B >
Answer #1

1) Comparative statement

Present Proposed
Sales 2000*90 = 180000 2000*1.1*90 = 198000
Variable expense 2000*63 = 126000 2000*1.1*65 = 143000
Fixed cost 30000 30000
Net operating income 24000 25000

Net operating income increase by $1000

2) Yes, Company should use high quality components